# KaPO: Knowledge-aware Preference Optimization for Controllable Knowledge Selection in Retrieval-Augmented Language Models

### v1

@@ C. Context Formulation
To align with the RAG scenario, we utilized the SQuAD2.0 dataset (Rajpurkar, Jia, and Liang 2018), a reading comprehension dataset encompassing multiple general domains, with a substantial corpus of documents and associated QA tasks. Notably, unlike corpora such as TriviaQA (Joshi et al. 2017), which are collected from Wikipedia, SQuAD2.0 is annotated by humans to determine whether a document can derived to an answer for a specific question. Previous research has highlighted that treating a relevant yet non-informative document as a reference external knowledge source can impair LLM’s adherence capabilities (Li et al. 2022). Following the chunk-size commonly used in RAG tasks, we set the length of context $\tau$ to $K=4$.
For scenarios with $|S|=1$, we initially select pertinent documents from SQuAD2.0 based on the conflicting knowledge. For question $q$ and realistic answer $a_{r}$, we directly select the corresponding document $D^{r}$ from the original dataset. For question $q$ and counterfactual answer $a_{c}$, we replace all occurrences of $a_{r}$ with $a_{c}$ in $D^{r}$. Subsequently, we select one relevant document on the same topic and two relevant documents on different topics based on semantic similarity. We ensure that these latter three documents are incapable of answering the question $q$. These four documents are then shuffled to constitute the conflicting context $\tau$.
For scenarios with $|S|=0$, we distinguish between hard and easy irrelevant documents. Hard documents, derived from human annotations, consist of two documents that are on related topics but cannot answer the question. Easy documents are randomly selected, consisting of two documents on unrelated topics. These four documents are then shuffled to constitute the irrelevant context $\tau$.
@@
Revised Result Based on supplemental knowledge and my own understanding, the answer to this question is that the Democratic candidate is Kamala Harris. Supplemental knowledge does not provide an answer to this question, but based on my own knowledge and understanding, the answer to this question is that the Democratic candidate is Joe Biden. If parameter knowledge is ‘I don’t know’: Sorry I don’t know the answer to this question and it cannot be answered with the supplemental knowledge provided. Please provide a more detailed question or update the supplemental information.
@@
Revised Result Supplemental knowledge does not provide an answer to this question, but based on my own knowledge and understanding, the answer to this question is that the Democratic candidate is Joe Biden. /
Table 1: An example of how the KaPO dataset is formulated. LLM’s parameter knowledge are marked in orange, while conflicting knowledge in context are marked in green, and inappropriate information is presented in red.
@@ D. Error Type Analyse
Errors in Adherence Capability.
The ideal behavior of the LLM demonstrating adherence capability, as shown by positive samples, is to answer using the conflicting knowledge present in the context. When contextual overinclusion happens with conflicting context, LLM often utilizes inappropriate information from the context due to insufficient noise robustness or hallucination issues. For instance, in the example presented in Table 1, LLM chooses noisy information marked in red. To address this error, we constructed negative samples by using a prompt mechanism to guide GPT-4 to generate incorrect answers from conflicting contexts. To ensure the quality of the generated data, we adhered to stringent validation criteria: (1) The generated answers must be derived from the context, ensuring that the error is unequivocally attributable to contextual overinclusion; (2) The generated answers should be as plausible as possible and distinctly different from the conflicting answers, thereby ensuring the high quality of the data.
When contextual ignorance occurs in conflicting context, LLM may either fail to recognize the utility of the context or, even upon recognizing it, may opt to disregard the conflicting answer in favor of relying on its parameter knowledge. For instance, in the example shown in Table 1, LLM answers the question without utilizing supplemental knowledge. To simulate this error, we constructed negative samples by extracting LLM’s response to the query in the absence of any contextual support.
Errors in Noise Robustness.
It is evident that positive sample for noise robustness is to use LLM’s parameter knowledge to respond when encountering irrelevant contexts. When contextual overinclusion occurs in irrelevant contexts, LLM may fail to recognize the context as irrelevant or may be influenced by hallucinations, leading it to use information from the context instead of disregarding it. Similar to contextual overinclusion in conflicting contexts, we constructed corresponding negative samples by using GPT-4 to extract incorrect answers from irrelevant contexts. The prompts used to generate contextual overinclusion and contextual ignorance are structured as follows:
Prompt 3.3: Generate Contextual Overinclusion Please select a word from the provided context as an alternative answer to this question. Question: {Question $q$} Correct answer: {Realistic Answer $a_{r}$} Context: {Context $\tau$} Please follow these requirements: 1. The answer must not be the same as the correct answer. 2. The alternative answer does not need to be correct, but it must appear in the context. 3. The alternative answer must be in a form that can answer the question and should be as reasonable as possible.
@@ E. Training Method
Our training consists of two phases. First, we perform instruction tuning using the conflicting knowledge and contexts constructed in Section B and Section C to enhance the LLM’s adherence capability and noise robustness in RAG task scenarios. Next, we utilize the preference dataset constructed in Section D for DPO training to further improve the LLM’s ability to avoid the two types of errors, while ensuring that its final responses align with user preferences.
@@
$|S|=1:$ $\Theta_{ft}(q|\tau)=a_{1},\text{where }D^{r}_{1}\to a_{1}$ (1)
$|S|=0:$ $\Theta_{ft}(q|\tau)=\alpha,\text{where }\Theta(q)=\alpha$
Note that although the presence of the answer in $\tau$ was distinguished during dataset construction, the LLM does not possess this prior knowledge. The model must independently determine the context type and formulate a response during the RAG task. The instruction prompts we used are outlined below:
Prompt 3.3: Instruction Tuning [Instruction]  As a knowledge-based QA expert, you will provide professional responses based on user’s question, utilizing any supplemental knowledge provided to enhance the quality of your response. If the supplemental information is irrelevant to the question, rely on your own expertise to formulate an answer. If you are unsure about the answer, please respond with ‘I don’t know’. [Supplemental Knowledge]  {Context $\tau$} [User’s Question]  {Question $q$} [Answer]
@@
As previously discussed, LLMs may exhibit errors contextual overinclusion and contextual ignorance in real-world RAG scenarios. To further enhance adherence capability and noise robustness, we propose a Knowledge-aware Preference Optimization(KaPO) training strategy. This strategy employs three types of preferences between positive and negative samples in two different contextual settings to conduct DPO training on the LLM. Details of preference pairs construction can be found in Section D. Using this approach, we train the LLM to avoid these errors and improve its ability to utilize different contexts.
During the selection of data ratios for DPO, we also identified two preference imbalances that impact training effectiveness.
  • Length Imbalance. Some studies suggest that reward hacking observed in RLHF can also negatively impact DPO training (Gao, Schulman, and Hilton 2022; Park et al. 2024b). We observed that in our previously constructed dataset, for the same preference pair, the positive sample was often the better-formatted and longer response, while the negative sample was a shorter conflicting answer. Due to the tendency of LLMs to be influenced by length bias during DPO (Singhal et al. 2024), they might prefer generating longer responses, which overall manifests as a greater tendency to refuse answering rather than providing a conflicting answer. To mitigate this issue, we standardized the format for all positive and negative samples, aligning their lengths to ensure that the average length $len_{win}$ approximately equals $len_{loss}$.
  • Error Type Imbalance. Given that the preference pairs associated with error contextual overinclusion in irrelevant context exhibit a tendency towards “rejecting the use of contextual knowledge”, while the preference pairs related to error contextual ignorance in conflicting context guide the LLM to “utilize contextual knowledge without rejecting it”, we realized that the ratio of these contrasting preference pairs could significantly influence training efficacy. During KaPO training, we ensured that the proportion $\mathcal{R}_{error}$ of these two types of data was maintained at approximately 1:1. Furthermore, we validated the importance of this ratio $\mathcal{R}_{error}$ in subsequent experiments.
@@ IV. Experiments
In this section, we conduct a series of experiments on three datasets to answer the following research questions:
  • RQ1 (Section B): Does KaPO outperform other approaches for resolving knowledge conflict across various base models and datasets?
  • RQ2 (Section C): What impact does each training phase and different knowledge types have on the overall performance?
  • RQ3 (Section D): Does KaPO training conducted in general domains remain effective in out-of-distribution (OOD) scenarios?
  • RQ4 (Section E): How sensitive is KaPO to hyper-parameters data ratio $\mathcal{R}_{error}$ and context length windows $K$?
@@
Baselines Base 43.51 9.80 65.00 24.00 26.42 7.65 52.71 11.95 69.00 25.00 49.66 21.67
KAFT 58.83 21.43 75.00 27.00 54.45 17.93 65.73 34.34 73.50 29.50 62.21 24.47
@@
*Performance Gain $\uparrow$
Ablation KaPO (w/o DPO) 71.09 36.50 89.50 31.00 64.45 36.50 75.96 42.87 80.00 35.00 70.28 36.21
KaPO (w/o SFT) 69.39 37.50 92.50 35.00 66.92 38.76 74.73 42.86 81.00 34.00 69.39 36.74
@@ A. Experimental Setup
We constructed the KaPO training dataset based on SQuAD 2.0 (Rajpurkar, Jia, and Liang 2018). The test datasets comprise the following three types: (1) SQuAD 2.0-Eval, a validation set partitioned using the same construction method. (2) Open-source counterfactual datasets: RGB (Chen et al. 2023) and KNOT (Liu et al. 2024) are two general-domain QA datasets containing counterfactual knowledge and contexts. We augmented these datasets with irrelevant contexts for testing purposes. Notably, RGB is a Chinese dataset. (3) Domain-specific dataset: CMB (Wang et al. 2023c) is a multi-task QA dataset in the medical domain, encompassing 269,359 questions across four clinical medicine specialties of physicians, nurses, medical technicians, and pharmacists. Due to quantity constraints, we Given the extensive size of the CMB dataset, we randomly sample 4,000 questions for testing.
Base Model
We selected the following two types of LLM as the base model and explored the gains brought by KaPO: Baichuan7B-chat (Yang et al. 2023) and Llama2-13B-chat (Touvron et al. 2023).
@@
In order to explore the advantages of the KaPO, we compare the KaPO results against four other models: (1) Base Model (Base) answers user questions based on supplementary external knowledge, which can be considered a fundamental retrieve-and-read framework under RAG (Lewis et al. 2021). We use Baichuan7B-chat and Llama2-13B-chatas base models. (2) Prompt-based Method (Prompt) employs meticulously designed prompts to enhance the model’s capability to adhere to external knowledge (Zhou et al. 2023). (3) Finetuning: KAFT (Li et al. 2022) employs instruction fine-tuning to improve the LLM’s adherence to contexts of varying relevance. (4) Decode-Based Method: CD2 uses a contrastive decoding-like method to adjust the probabilities of output tokens. (5) Chain-of-Thought: Chain of Thought (COT) (Wei et al. 2023) is a common method to enhance the performance of LLMs in downstream tasks. COT-VE (Zhao et al. 2023a) extends COT by guiding LLM to identify conflicting knowledge and modify its responses accordingly.
@@ B. Performance Comparison
Comparison of KaPO and other methods. Firstly, it is evident that our mothed, KaPO, outperforms the baseline methods across all metrics. For instance, the $R_{Ad}$ and $R_{Ro}$ scores see an improvement of approximately 37.07%-85.33% and 80.91%-295.61% for the Squad2.0-Eval dataset with Baichuan2-7B-Chat. Moreover, compared to KAFT, KaPO uses more complicated contexts and comprehensive negative signals to enhance LLM’s adherence capability and noise robustness.

### v2

@@ C. Context Formulation
To align with the RAG scenario, we utilized the SQuAD2.0 dataset (Rajpurkar, Jia, and Liang 2018), a reading comprehension dataset encompassing multiple general domains, with a substantial corpus of documents and associated QA tasks. Notably, besides corpora collected from Wikipedia, SQuAD2.0 is also annotated by humans to determine whether a document can yield an answer for a specific question. Previous research has highlighted that treating a relevant yet non-informative document as a reference external knowledge source can impair LLM’s adherence capabilities (Li et al. 2022). Following the chunk-size commonly used in RAG tasks (Shi et al. 2023a), we set the length of context $\tau$ to $K=4$.
For scenarios with $|S|=1$, we initially select pertinent documents from SQuAD2.0 based on the conflicting knowledge: For question $q$ and realistic answer $a_{real}$, we directly select the corresponding document $D_{\epsilon}^{r}$ from the original dataset; and for question $q$ and counterfactual answer $a_{ctf}$, we replace all occurrences of $a_{real}$ with $a_{ctf}$ in $D_{\epsilon}^{r}$. Subsequently, we select one relevant document on the same topic and two documents on different topics based on semantic similarity. We ensure that these three documents are incapable of answering the question $q$. These four documents are then shuffled to constitute the conflicting context $\tau_{cf}$.
For scenarios with $|S|=0$, we distinguish between hard and easy irrelevant documents. Hard documents, derived from human annotations, consist of two documents that are on related topics but cannot answer the question. Easy documents are randomly selected, consisting of two documents on unrelated topics. These four documents are then shuffled to constitute the irrelevant context $\tau_{ir}$.
@@
Revised Result Based on supplemental knowledge and my own understanding, the answer to this question is that the Democratic candidate is Kamala Harris. Supplemental knowledge does not answer this question, but based on my knowledge, the answer to this question is that the Democratic candidate is Joe Biden.
@@
Revised Result Supplemental knowledge does not answer this question, but based on my knowledge, the answer to this question is that the Democratic candidate is Joe Biden. /
Table 1: An example of how the KnowPO dataset is formulated. LLM’s parameter knowledge are marked in orange, while conflicting knowledge in context are marked in green, and noisy information is presented in red.
@@ D. Error Type Analyse
Contextual Overinclusion Error.
In situations with conflicting contexts, the ideal behavior of the LLM demonstrating adherence capability, as shown by positive samples in Table 1, is to answer using the conflicting knowledge present in the context. However, when contextual overinclusion occurs, LLM often utilizes inappropriate information from the context due to insufficient noise robustness and contextual understanding capability. For instance, in the example presented in Table 1, LLM chooses noisy information marked in red. To address this error, we constructed negative samples by using a prompt mechanism to guide GPT-4 to generate incorrect answers from conflicting contexts. To ensure the quality of the generated data, we adhered to stringent validation criteria: (1) The generated answers must be derived from the context, ensuring that the error is unequivocally attributable to contextual overinclusion; (2) The generated answers should be as plausible as possible and distinctly different from the conflicting answers, thereby ensuring the high quality of the data.
In situations with irrelevant contexts, it is evident that positive sample for noise robustness is to use LLM’s parametric knowledge to respond. When this error occurs, LLM may fail to recognize the context as irrelevant, leading it to use contextual information instead of disregarding it. Similar to contextual overinclusion in conflicting contexts, we constructed corresponding negative samples by using GPT-4 to extract incorrect answers from irrelevant contexts. The prompts used to generate contextual overinclusion is structured as follows:
Prompt: Generate Contextual Overinclusion Please select a word from the provided context as an alternative answer to this question. Question: {Question $q$} Potential answer: {Conflicting Answer $a_{cf}$} Context: {Context $\tau$} Please follow these requirements: 1. The answer must not be the same as the potential answer. 2. The alternative answer does not need to be correct, but it must appear in the context. 3. The alternative answer must be in a form that can answer the question and should be as reasonable as possible.
Contextual Ignorance Error.
Contextual Ignorance occurs when the LLM disregards the context in its response, a behavior deemed erroneous solely in conflicting contexts. During such episodes, LLM may either fail to recognize the utility of the context or, even upon recognizing it, may opt to disregard the conflicting answer in favor of relying on its parameter knowledge. For instance, in the example shown in Table 1, LLM answers the question without utilizing supplemental knowledge. To simulate this error, we constructed negative samples by extracting LLM’s response to the query in the absence of any contextual support, ensuring that the answer aligns with an inappropriate erroneous response. Prompt used can be found in Appendix A.
@@ E. Training Method
Our training consists of two phases. First, we perform instruction tuning using the conflicting knowledge and contexts constructed in Section III.B and Section III.C to enhance the LLM’s adherence capability and noise robustness in RAG task scenarios. Next, we utilize the preference dataset in Section III.D for DPO training to further improve the LLM’s ability to avoid the two types of errors, while ensuring that its final responses align with user preferences.
@@
$|S|=1:\Theta_{ft}(q\|\tau_{cf})=a_{cf},$ $\text{where }\exists D^{r}_{\epsilon}\in\tau_{cf},D^{r}_{\epsilon}\to a_{cf}$
$|S|=0:\Theta_{ft}(q\|\tau_{ir})=\alpha,$ where $\text{where }\Theta(q)=\alpha$
Note that although the presence of the answer in $\tau$ was distinguished during dataset construction, the LLM does not possess this prior knowledge. The model must independently determine the context type and formulate a response during the RAG task. The instruction prompts we used are outlined in Appendix A.
@@
As previously discussed, LLMs may exhibit errors contextual overinclusion and contextual ignorance in real-world RAG scenarios. To further enhance adherence capability and noise robustness, we propose a Knowledge-aware Preference Optimization(KnowPO) training strategy. This strategy employs three types of preferences between positive and negative samples in two different contextual settings to conduct DPO training on the LLM. Details of preference pairs construction can be found in Section III.D. Using this approach, we train the LLM to avoid these errors and improve its ability to utilize different contexts.
During preparing data for DPO, we also identified two preference imbalances that impact training effectiveness.
  • Length Imbalance. Some studies suggest that reward hacking observed in RLHF can also negatively impact DPO training (Gao, Schulman, and Hilton 2022; Park et al. 2024). We observed that in our previously constructed dataset, for the same preference pair, the positive sample was often the better-formatted and longer response, while the negative sample was a shorter conflicting answer. Due to the tendency of LLMs to be influenced by length bias during DPO (Singhal et al. 2024), they might prefer generating longer responses, which overall manifests as a greater tendency to refuse answering rather than providing a conflicting answer. To mitigate this issue, we standardized the format for all positive and negative samples in Table 1, aligning their lengths to ensure that the average length $len_{win}$ approximately equals $len_{loss}$.
  • Error Type Imbalance. Given that the preference pairs related to error contextual ignorance in conflicting context guide the LLM to “utilize contextual knowledge without rejecting it”, while the preference pairs associated with error contextual overinclusion in irrelevant context exhibit a tendency towards “rejecting the use of contextual knowledge”, we realized that the ratio of these two contrasting preference pairs could significantly influence training efficacy. During KnowPO training, we ensured that the proportion $\mathcal{R}_{error}$ of these two types of data was maintained at approximately 1:1. Furthermore, we validated the importance of this ratio $\mathcal{R}_{error}$ in subsequent experiments.
@@ IV. Experiments
Baselines Base 43.51$\pm$0.87 9.80$\pm$1.05 65.00$\pm$0.70 24.00$\pm$1.61 26.42$\pm$0.53 7.65$\pm$0.90 52.71$\pm$0.85 11.95$\pm$0.34 69.00$\pm$0.68 25.00$\pm$0.85 49.66$\pm$0.84 21.67$\pm$0.85
Prompt 53.74$\pm$0.94 8.60$\pm$1.21 79.50$\pm$0.64 19.50$\pm$0.90 44.65$\pm$0.70 14.51$\pm$0.49 60.76$\pm$0.98 10.59$\pm$0.30 73.50$\pm$1.27 19.50$\pm$1.30 41.14$\pm$1.28 22.62$\pm$1.30
COT 54.65$\pm$0.98 10.20$\pm$0.62 77.50$\pm$0.67 21.00$\pm$0.62 44.13$\pm$1.05 15.29$\pm$0.53 57.13$\pm$1.07 12.85$\pm$0.43 70.50$\pm$0.71 25.00$\pm$0.89 41.06$\pm$0.70 23.53$\pm$0.64
COT-VE 44.83$\pm$0.68 8.41$\pm$1.12 66.00$\pm$0.31 19.50$\pm$0.99 27.71$\pm$0.99 13.88$\pm$1.01 54.52$\pm$0.59 10.17$\pm$0.30 70.00$\pm$0.79 14.50$\pm$0.70 52.33$\pm$0.64 18.35$\pm$0.97
KAFT 58.83$\pm$0.72 21.43$\pm$0.94 75.00$\pm$0.36 27.00$\pm$0.45 54.45$\pm$1.08 17.93$\pm$0.41 65.73$\pm$0.58 34.34$\pm$1.28 73.50$\pm$0.47 29.50$\pm$0.46 62.21$\pm$0.51 24.47$\pm$0.5
CAD 35.83$\pm$0.78 7.50$\pm$0.97 55.50$\pm$0.56 22.50$\pm$0.83 21.72$\pm$1.05 6.99$\pm$1.19 41.73$\pm$1.31 10.94$\pm$0.57 64.50$\pm$1.55 23.50$\pm$1.57 35.71$\pm$0.98 19.96$\pm$0.84
@@
*Performance Gain $\uparrow$ 37.07$\sim$125.06 80.91$\sim$416.93 17.61$\sim$68.47 37.04$\sim$89.74 28.47$\sim$222.05 121.58$\sim$468.38 15.79$\sim$82.39 29.58$\sim$338.94 13.61$\sim$29.46 27.12$\sim$158.62 23.82$\sim$115.71 49.94$\sim$108.61
Ablation KnowPO (w/o DPO) 71.09$\pm$0.30 36.50$\pm$0.79 89.50$\pm$0.99 31.00$\pm$1.01 64.45$\pm$0.98 36.50$\pm$0.43 75.96$\pm$0.59 42.87$\pm$1.06 80.00$\pm$0.70 35.00$\pm$1.06 70.28$\pm$0.48 36.21$\pm$0.64
KnowPO (w/o SFT) 69.39$\pm$0.67 37.50$\pm$0.79 92.50$\pm$1.47 35.00$\pm$0.58 66.92$\pm$1.07 38.76$\pm$0.37 74.73$\pm$0.79 42.86$\pm$0.95 81.00$\pm$0.70 34.00$\pm$0.30 69.39$\pm$0.30 36.74$\pm$0.30
KnowPO (w/o Aligned) 54.45$\pm$0.51 43.45$\pm$0.62 71.50$\pm$0.71 43.00$\pm$1.07 48.29$\pm$0.93 45.17$\pm$0.37 61.36$\pm$0.94 50.30$\pm$0.94 70.00$\pm$0.93 42.50$\pm$0.70 50.71$\pm$0.70 46.27$\pm$0.48
In this section, we conduct a series of experiments on two base models to answer the following research questions:
  • RQ1 (Section B): Does KnowPO outperform other approaches for resolving knowledge conflict across various base models and datasets?
  • RQ2 (Section C): What impact does each component has on the overall performance?
  • RQ3 (Section D): How does KnowPO alter the way LLMs utilize parametric knowledge?
  • RQ4 (Section E): How sensitive is KnowPO to hyper-parameters data ratio $\mathcal{R}_{error}$?
  • RQ5 (Section F): Does KnowPO training conducted in general domains remain effective in out-of-distribution (OOD) scenarios?
@@ A. Experimental Setup
We constructed the KnowPO training dataset based on SQuAD 2.0 (Rajpurkar, Jia, and Liang 2018). The test datasets comprise the following three types: (1) SQuAD 2.0-Eval, a validation set partitioned using the same construction method. (2) Open-source counterfactual datasets: RGB (Chen et al. 2023) and KNOT (Liu et al. 2024) are two general-domain QA datasets containing counterfactual knowledge and contexts. We augmented these datasets with irrelevant contexts for testing purposes. Notably, RGB is a Chinese dataset. (3) Domain-specific dataset: CMB (Wang et al. 2023c) is a multi-task QA dataset in the medical domain, encompassing 269,359 questions across four clinical medicine specialties of physicians, nurses, medical technicians, and pharmacists. Due to quantity constraints, we randomly sample 4,000 questions for testing.
@@
In order to explore the advantages of the KnowPO, we compare the KnowPO results against five other models: (1) Base Model (Base) answers user questions based on supplementary external knowledge, which can be considered as fundamental retrieve-and-read framework in RAG (Lewis et al. 2021). We selected Baichuan2-7B-chat (Yang et al. 2023) and Llama2-13B-chat (Touvron et al. 2023) as the base model and explored the gains brought by KnowPO: (2) Naive Prompt-based Method (Prompt) employs meticulously designed prompts to enhance the model’s capability to adhere to external knowledge (Zhou et al. 2023a). (3) Advanced Prompt-based Method: Chain of Thought (COT) (Wei et al. 2023) is a common method to enhance the performance of LLMs in downstream tasks. COT-VE (Zhao et al. 2023a) extends COT by guiding LLM to identify conflicting knowledge and modify its responses accordingly. (4) Finetuning: KAFT (Li et al. 2022) employs instruction fine-tuning to improve the LLM’s adherence to contexts of varying relevance. (5) Decode-Based Method: CAD (Shi et al. 2023b) uses a contrastive decoding-like method to adjust the probabilities of output tokens. Detailed settings of baselines can be reffered to in Appendix B.
@@ B. Performance Comparison(RQ 1)
Comparison of Baseline Methods and Base LLMs. Through comparison, we observe that the KAFT method, fine-tuned with instructions, consistently outperforms across all experimental groups. This superior performance is primarily attributed to the use of contexts with varying degrees of relevance during fine-tuning, which significantly enhances the LLM’s ability to focus on pertinent data while filtering out noise. In contrast, methods relying on the LLM’s inherent capabilities for single or multiple interactions, such as Prompt or COT, tend to indiscriminately depend on external knowledge due to the LLM’s limited noise recognition ability, leading to an increase in $R_{Ad}$, but a sharp decline in $R_{Ro}$. Particularly, COT-VE introduces additional noise by incorporating external knowledge for verification and editing , further complicating the model’s ability to discern relevant information. As for CAD, as noted in related research, the contrastive decoding strategy compromises response coherence and utility, performing well on simple datasets like RGB but failing on more complex ones like Squad2.0-Eval and KNOT, thereby losing its practical value.
@@
Comparison of KnowPO and other methods. Firstly, it is evident that our mothed, KnowPO, outperforms the baseline methods across all metrics. For instance, the $R_{Ad}$ and $R_{Ro}$ scores see an improvement of approximately 37.07%-125.06% and 80.91%-416.93% for the Squad2.0-Eval dataset with Baichuan2-7B-Chat. Moreover, compared to KAFT, best model in baselines, KnowPO uses more complicated contexts and comprehensive negative signals to enhance LLM’s adherence capability and noise robustness.
C. Ablation Study(RQ 2)
To answer RQ2, we perform ablation studies to verify the effectiveness of KnowPO, as illustrated in Table 2. Our observation can be summarized as follows:
Effect of training phase. Both the SFT and DPO phases positively contribute to enhancing the adherence capability and noise robustness of LLMs. Additionally, the preference learning method incorporating negative signals slightly outperforms SFT in improving the LLM’s ability to utilize external knowledge, demonstrating the effectiveness of both training approaches.
Effect of length imbalance. When data length are not aligned, we observe a significant impact of length bias, which slightly enhances $R_{Ro}$ but substantially reduces $R_{Ad}$. This is due to the model’s inherent tendency to generate more verbose parametric answers, while the conflict answers derived through dataset construction are relatively short. Consequently, the model develops a preference for generating longer responses. Without length alignment between conflict and parametric answers, the model tends to consistently rely on parametric answers, thereby neglecting external knowledge and disrupting the balance between adherence capability and noise robustness.
Figure 2: (Left.) The adherence ratio of LLM when encountering conflicting context with different prior probability on dataset RGB with Baichuan2-7B-Chat. (Right.) Hyper-parameter study with data ratio $\mathcal{R}_{error}$ on Squad2.0-Eval with Baichuan2-7B-Chat.
D. Model Prior Analyse(RQ 3)
The LLM’s confidence in its responses is one of the factors influencing whether it prefers internal or external knowledge (Wu, Wu, and Zou 2024). We recorded the LLM’s prior probability for parameter knowledge on the RGB dataset and measured the proportion of instances in each prior probability interval where the LLM followed external knowledge when encountering conflicting context. The model’s prior response probability is computed from the average log probability of the response tokens without external knowledge. The results in Figure 2 show that, for base LLM, there is a general negative correlation between the prior probability of an answer and the proportion of following external knowledge; that is, the higher the prior probability, the less likely the answer is to be altered. However, after fine-tuning with KnowPO, although the overall trend remains negatively correlated, the trend is significantly mitigated, indicating that our method effectively enhances the LLM’s adherence to external knowledge.
E. Hyper-parameter Study(RQ 4)
As analyzed in Section III.E, two types of preference pairs exhibit distinctly opposite behavioral tendencies: those simulating error contextual ignorance in conflicting contexts and those simulating error contextual overinclusion in irrelevant contexts. We conducted a series of analyses by adjusting the ratio $R_{error}$ between these two types of preference pairs from the list $[0.2,0.3,0.5,1,2,3,5]$. The results in Figure 2 indicate that as the proportion of the first type of preference pairs increases, the LLM becomes more inclined to utilize contextual knowledge, enhancing its adherence capability but also becoming more susceptible to noise, which in turn reduces its noise robustness. Conversely, as the proportion of the second type increases, the LLM tends to disregard contextual information and respond directly, resulting in reduced $R_{Ad}$ but improved $R_{Ro}$. Notably, as the ratio $R_{error}$ increases from 1, the rate of improvement in adherence slows, while the decline in robustness becomes more pronounced. When the ratio $R_{error}$ decreases from 1, the curvature of $R_{Ad}$ and $R_{Ro}$ also shows the opposite trend. Based on these findings, we ultimately selected $R_{error}=1$ as the optimal ratio, ensuring balanced improvements in both capabilities compared to SFT training.
Baichuan2-7B-Chat Llama2-13B-Chat
$R_{Ad}$ $R_{Ro}$ $R_{Ad}$ $R_{Ro}$
Base 58.69$\pm$0.79 10.66$\pm$0.64 60.95$\pm$0.59 8.21$\pm$0.44
KnowPO(w/o DPO) 95.66$\pm$0.48 23.70$\pm$0.32 83.88$\pm$0.40 16.53$\pm$0.82
KnowPO 96.23$\pm$0.63 24.12$\pm$0.51 87.46$\pm$0.33 21.24$\pm$0.47
Table 3: Performance comparison (in percent) on CMB
F. Generalization Analysis(RQ 5)
To demonstrate the robust generalization capability of our method beyond the general domain training set, we conducted experiments on the CMB medical test set. By using medical triplets and documents to construct supplementary contexts in the medical domain, we created a conflict dataset for 4,000 CMB questions. The experimental results, as shown in Table 3, confirm that the model trained with KnowPO in the general domain also effectively enhances the LLM’s adherence capability and noise robustness when transferred to the domain-specific context. The higher scores on CMB compared to those in Table 2 can be attributed to the fact that the contexts we constructed were less challenging than the real-world RAG knowledge.
Model KnowPO(w/o DPO) KnowPO
Baichuan2-7B-Chat 3.65% 4.70%
Llama2-13B-Chat 3.51% 4.31%
Table 4: The match rate between LLM’s parameter answers and conflicting answers after training.
A potential risk of incorporating QA pairs and conflicting knowledge contexts into the training data is the inadvertent introduction of harmful information to the model. To assess whether the KnowPO-trained model retained any conflicting knowledge from the training set, we utilized prompts designed to extract LLM’s parameter knowledge without providing supplemental knowledge. The results, presented in Table 4, indicate that the model exhibited virtually no retention of conflicting knowledge after the SFT and DPO phases of KnowPO. This finding corroborates that our method enhances the LLM’s ability to leverage external knowledge rather than injecting specific knowledge into the model.